# Against All Odds: Overcoming Typology, Script, and Language Confusion in Multilingual Embedding Inversion Attacks

### v1

@@ Abstract
Large Language Models (LLMs) are susceptible to malicious influence by cyber attackers through intrusions such as adversarial, backdoor, and embedding inversion attacks. In response, the burgeoning field of LLM Security aims to study and defend against such threats. Thus far, the majority of works in this area have focused on monolingual English models, however, emerging research suggests that multilingual LLMs may be more vulnerable to various attacks than their monolingual counterparts. While previous work has investigated embedding inversion over a small subset of European languages, it is challenging to extrapolate these findings to languages from different linguistic families and with differing scripts. To this end, we explore the security of multilingual LLMs in the context of embedding inversion attacks and investigate cross-lingual and cross-script inversion across 20 languages, spanning over 8 language families and 12 scripts. Our findings indicate that languages written in Arabic script and Cyrillic script are particularly vulnerable to embedding inversion, as are languages within the Indo-Aryan language family. We further observe that inversion models tend to suffer from language confusion, sometimes greatly reducing the efficacy of an attack. Accordingly, we systematically explore this bottleneck for inversion models, uncovering predictable patterns which could be leveraged by attackers. Ultimately, this study aims to further the field’s understanding of the outstanding security vulnerabilities facing multilingual LLMs and raise awareness for the languages most at risk of negative impact from these attacks.
@@ Introduction
Vector databases store texts encoded into embeddings, and represent a 1.5 billion USD market globally as of 2023, which is expected to grow to an impressive 4 billion within the next 5 years (Markets and Markets 2023). The security of this technology is increasingly critical for providers and customers depending on vector databases for their businesses. Presently, vector databases are known to have one outstanding security vulnerability: text embeddings can be reverted into original text, even without prior knowledge of the embedding model, thus constituting a major threat to privacy (Song and Raghunathan 2020; Höhmann, Rettinger, and Kugler 2021; Hayet, Yao, and Luo 2022; Morris et al. 2023; Li, Xu, and Song 2023; Chen, Lent, and Bjerva 2024). In other words, text embeddings are no more secure than plain text (Morris et al. 2023). A malicious actor needs only to intercept the vectors in order to carry out an embedding inversion attack, whereby potentially sensitive information is leaked.
@@
As with most NLP problems, large English datasets have served as the primary focus of research on embedding inversion attacks. Emerging work, however, highlights how the field’s over-reliance on English comes at a grave cost to the general security of Large Language Models (LLMs). For example, Yong, Menghini, and Bach (2024) demonstrate how ChatGPT’s built-in safety features can be easily by-passed by prompting in lower-resourced languages, while OpenAI’s defences function as expected for English and other highly-resourced languages. Meanwhile, Wang et al. (2024a) show how low-resource languages can be weaponized to poison models in backdoor attacks against machine translation systems. In the context of embedding inversion attacks, security may also be compromised by failing to consider a diversity of languages, especially as vector database providers may leverage multilingual LLMs to serve clients around the globe. Chen, Lent, and Bjerva (2024) find that multilingual LLMs are more vulnerable to embedding inversion attacks than their monolingual counterparts. However, their study includes only four highly-resourced European languages, restricted to Latin script. Conclusions drawn over a small set of relatively homogeneous languages cannot be trivially extended to massively multilingual LLMs, thus stressing the need for investigations of embedding inversion attacks with languages of diverse linguistic families and varying scripts.
@@
In this work, we explore the nuances of embedding inversion attacks against multilingual models across 20 languages from 8 language families written in 12 scripts. We find that the success of inversion attacks is largely influenced by combinations of language family and script, ultimately showing that some languages are more vulnerable to attacks than others. Beyond elucidating threats facing multilingual encoders for vector databases, we analyze a common hindrance to the performance of embedding inversion: language confusion, a term recently coined by Marchisio et al. (2024), defined as a limitation of LLMs where the models are often unable to consistently generate text in the user’s desired language, or the appropriate language given the context. We show that language confusion can be predicted and controlled for, unfortunately to the benefit of would-be attackers.
@@
    We investigate embedding inversion attacks across 8 language families, comparing within and across-family and find that the Indo-Aryan family was the most vulnerable.
@@
    We examine attacks across 12 scripts, finding that scripts differ in vulnerability to inversion attacks. Our findings indicate that Arabic and Cyrillic scripts were the most susceptible to attack.
@@
Language Confusion Language confusion is a significant limitation of LLMs, where a model cannot consistently generate text in the desired language or appropriate language given the context (Marchisio et al. 2024). This can result in full-response confusion, line-level confusion, and word-level confusion (Vu et al. 2022; Marchisio et al. 2024). In embedding inversion attacks, Chen, Lent, and Bjerva (2024) attributes language confusion to using a multilingual black-box encoder. However, in this work, we find that language confusion can occur when with both a multilingual and monolingual black-box encoder. It should be noted that language confusion is unlikely to directly thwart attackers. For example, machine translation can transform uncovered text to another language. Still, as we find that language confusion incurs some cost for attack models, regarding accuracy, investigation is warranted to ascertain the threat level to massively multilingual LLMs.
@@
In this work, we consider a scenario where a malicious attacker has illegitimate access to stolen embeddings and API access to the black-box encoder. To gauge the vulnerability of multilingual models against black-box inversion embedding attacks, we implement inversion attacks on a diverse setting of languages across language families and scripts. The attack scenario is formally defined as follows: given a text sequence $x$ and a black-box encoder $\phi$, an external inversion model $\psi$ aims to recover the text $x$ from its embeddings $\phi(x)$. The objective can be formalized as such that:
@@
This forms a text-to-text generation task, resulting in the trained base model. However, this approach alone is insufficient. To improve the inversion performance, a corrector model is implemented, where $\phi$ is queried using the output of the base model, and new hypothesis embeddings $\hat{e}^{(t)}=\phi(x^{(t)})$ is computed in iteration $t$ with which a new correction text $x^{(t+1)}$ is generated. The model is defined by marginalizing over intermediate hypotheses (Morris et al. 2023):
@@
Attacking In-script Textual Embeddings
@@
Attacking In-Family Text Embeddings
@@ Conclusion
Our work highlights the fact that vulnerabilities in multilingual embedding attacks disproportionately affect low-resource languages. This risk is exacerbated by the fact that language confusion, which can be exploited by attackers, is shown to be both predictable and solvalbe. Ultimately, the security of LLMs and guarantees for safety for all end-users relies on the inclusion of diverse languages in LLM security research. Multilinguality poses unique challenges, and solutions offered for high-resource languages are not necessarily useful for those with fewer resources. Without prioritizing diversity in this emerging field, LLMs will remain vulnerable, with marginalized communities worldwide facing disproportionate risk as LLMs become more widespread.
Open Source
We make all our code, used datasets and trained inversion models publicly available in GitHub https://github.com/siebeniris/vec2text˙exp/tree/aaai and Hugging Face https://huggingface.co/yiyic.
@@
A core limitation of this work is the computationally intense experiments, requiring approximately 40,000 GPU computing hours. While expanding this research direction to more languages will further increase this expense, we advocate for ensuring that languages other than English are not left behind in terms of NLP security.
@@
This work investigates attacks on multilingual embedding models and the phenomenon of language confusion, with the goal of highlighting vulnerabilities across languages representing diverse scripts and typologies. We encourage the community to incorporate a broader range of languages in NLP security research. Some of the languages we examine, such as Yiddish and Meadow Mari, are particularly vulnerable. To minimize potential harm, we do not train any attacks on these languages, instead focusing solely on evaluation to determine their current threat level. The language models used in this study are open-source, ensuring that this research does not pose an immediate threat to EaaS providers, who are likely using private models. Additionally, we take care not to experiment with any sensitive data, ensuring that no real-world harm results from this work.
@@ Acknowledgements
YC, HCL, and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454). We further acknowledge the support of the AAU AI Cloud and express our gratitude to DeiC for providing computing resources on the LUMI cluster (project no. DeiC-AAU-N5-2024085-H2-2024-28). We extend our thanks to Sighvatur Sveinn Davidsson and Frederik Petri Svenningsen for their assistance with AI Cloud, and to Esther Ploeger and Mike Zhang for their valuable discussions related to this work.

### v2

@@ Abstract
Large Language Models (LLMs) are susceptible to malicious influence by cyber attackers through intrusions such as adversarial, backdoor, and embedding inversion attacks. In response, the burgeoning field of LLM Security aims to study and defend against such threats. Thus far, the majority of works in this area have focused on monolingual English models, however, emerging research suggests that multilingual LLMs may be more vulnerable to various attacks than their monolingual counterparts. While previous work has investigated embedding inversion over a small subset of European languages, it is challenging to extrapolate these findings to languages from different linguistic families and with differing scripts. To this end, we explore the security of multilingual LLMs in the context of embedding inversion attacks and investigate cross-lingual and cross-script inversion across 20 languages, spanning over 8 language families and 12 scripts. Our findings indicate that languages written in Arabic script and Cyrillic script are particularly vulnerable to embedding inversion, as are languages within the Indo-Aryan language family. We further observe that inversion models tend to suffer from language confusion, sometimes greatly reducing the efficacy of an attack. Accordingly, we systematically explore this bottleneck for inversion models, uncovering predictable patterns which could be leveraged by attackers. Ultimately, this study aims to further the field’s understanding of the outstanding security vulnerabilities facing multilingual LLMs and raise awareness for the languages most at risk of negative impact from these attacks. [^1] GitHub: https://github.com/siebeniris/vec2text˙exp, HuggingFace: https://huggingface.co/yiyic.
@@ Introduction
Vector databases store texts encoded into embeddings, representing a 1.5 billion USD market globally as of 2023, which is expected to grow to an impressive 4 billion within the next 5 years (Markets and Markets 2023). In more concrete contexts, services like Pinecone and Qdrant, [^2] https://www.pinecone.io/ and https://qdrant.tech/ store embeddings for retrieval in applications like retrieval-augmented generation (RAG) (Lewis et al. 2020). In RAG, user queries are embedded via LLMs and sent to a vector database for semantic search, while large-scale embeddings of text data are also stored for semantic search. The security of this technology is increasingly critical for providers and customers depending on vector databases for their businesses. Presently, vector databases are known to have one outstanding security vulnerability: text embeddings can be reverted into original text, even without prior knowledge of the embedding model, thus constituting a major threat to privacy (Song and Raghunathan 2020; Höhmann, Rettinger, and Kugler 2021; Hayet, Yao, and Luo 2022; Morris et al. 2023; Li, Xu, and Song 2023; Chen, Lent, and Bjerva 2024). In other words, text embeddings are no more secure than plain text (Morris et al. 2023). A malicious actor needs only to intercept the vectors in order to carry out an embedding inversion attack, whereby potentially sensitive information is leaked.
@@
As with most NLP problems, large English datasets have served as the primary focus of research on embedding inversion attacks. Emerging work, however, highlights how the field’s over-reliance on English comes at a grave cost to the general security of Large Language Models (LLMs). For example, Yong, Menghini, and Bach (2024) demonstrate how ChatGPT’s built-in safety features can be easily by-passed by prompting in lower-resourced languages, while OpenAI’s defences function as expected for English and other highly-resourced languages. Meanwhile, Wang et al. (2024a) show how low-resource languages can be weaponized to poison models in backdoor attacks against machine translation systems. In the context of embedding inversion attacks, security may also be compromised by failing to consider a diverse range of languages, especially as vector database providers may leverage multilingual LLMs to serve clients around the globe. Chen, Lent, and Bjerva (2024) find that multilingual LLMs are more vulnerable to embedding inversion attacks than their monolingual counterparts. However, their study includes only four highly-resourced European languages, restricted to Latin script. Conclusions drawn over a small set of relatively homogeneous languages cannot be trivially extended to massively multilingual LLMs, thus stressing the need for investigations of embedding inversion attacks with languages of diverse linguistic families and varying scripts.
@@
In this work, we explore the nuances of embedding inversion attacks against multilingual models across 20 languages from 8 language families written in 12 scripts. We find that the success of inversion attacks is largely influenced by combinations of language family and script, ultimately showing that some languages are more vulnerable to attacks than others. Beyond elucidating threats facing multilingual encoders for vector databases, we analyze a common hindrance to the performance of embedding inversion: language confusion, a term recently coined by Marchisio et al. (2024), defined as a limitation of LLMs where the models are often unable to consistently generate text in the user’s desired language, or the appropriate language given the context. Fig. 1 illustrates different vulnerability levels in inversion attacks and language confusion. Specifically, Arabic textual embeddings (input) are inverted by a model trained on German embeddings, with the final output producing a mostly coherent German sentence containing an English term. Language confusion is understood as an “erroneous behavior” across the whole of NLP, reported in LLM instruction tuning and is detrimental to generation performance in terms of intelligibility in general (Marchisio et al. 2024). We show that language confusion can be predicted and controlled for, unfortunately to the benefit of would-be attackers.
@@
    We investigate embedding inversion attacks across 8 language families, comparing within and across-family and find that the Indo-Aryan family was the most vulnerable. Notably, inverting Punjabi texts produces meaningful sentences only in the in-family training setting compared to training it independently.
@@
    We examine attacks across 12 scripts, finding that scripts differ in vulnerability to inversion attacks. Our findings indicate that Arabic and Cyrillic scripts were the most susceptible to attack. In-script training boosts Urdu inversion attack performance, doubling the BLEU score.
@@
Language Confusion Language confusion is a significant limitation of LLMs, where a model cannot consistently generate text in the desired language or appropriate language given the context (Marchisio et al. 2024). This can result in full-response confusion, line-level confusion, and word-level confusion (Vu et al. 2022; Marchisio et al. 2024). In embedding inversion attacks, Chen, Lent, and Bjerva (2024) attributes language confusion to using a multilingual black-box encoder. However, in this work, we find that language confusion can occur when with both a multilingual and monolingual black-box encoder. Language confusion as a vulnerability is also supported by recent work, whereby it has already been proven an effective weapon for jailbreaking state-of-the-art LLMs (Song et al. 2024).
@@
In this work, we consider a scenario where a malicious attacker has illegitimate access to stolen embeddings and API access to the black-box encoder. To gauge the vulnerability of multilingual models against black-box inversion embedding attacks, we implement inversion attacks on a diverse setting of languages across language families and scripts. The attack scenario is formally defined as follows: given a text sequence $x$ from a dataset $D$, and a black-box encoder $\phi$, an external inversion model $\psi$ aims to recover the text $x$ from its embeddings $\phi(x)$. The objective can be formalized as such that:
@@
This forms a text-to-text generation task, resulting in the trained base model. However, this approach alone is insufficient. To improve the inversion performance, a corrector model is implemented, where $\phi$ is queried using the output of the base model, and new hypothesis embeddings $\hat{e}^{(t)}=\phi(x^{(t)})$ is computed in iteration $t$ with which a new correction text $x^{(t+1)}$ is generated. The corrector aims to find text $\hat{x}$ that maximizes cosine similarity with the target embedding $\phi(x)$ : $\hat{x}=\arg\max_{x}\cos(\phi(x),e)$. At step $(t+1)$, the model concatenates the previous output $\hat{x}^{(t)}$, hypothesis embeddings $\hat{e}^{(t)}$, and target embeddings $\hat{e}^{t}$ as in the equation  (3) in the methodology section. The model is defined by marginalizing over intermediate hypotheses (Morris et al. 2023):
@@
Monolingual vs. Multilingual LLMs
@@
Attacking In-script Textual Embeddings
@@ Conclusion
Our work highlights the fact that vulnerabilities in multilingual embedding attacks disproportionately affect low-resource languages. This risk is further exacerbated by the fact that language confusion, which can be exploited by attackers, is shown to be both predictable and solvable. We prove this by showing directly that attackers drastically increase the efficacy of their embedding inversion attacks, when specifically accounting for language confusion (i.e., by training on in-family and in-script languages), making it an appealing exploit for malicious actors to potentially leverage. In this context, we argue that a deeper understanding of language confusion, including its connection to linguistic characteristics, could offer critical insights into the mechanics of multilingual attacks. Ultimately, the security of LLMs and guarantees for safety for all end-users relies on the inclusion of diverse languages in LLM security research. Multilinguality poses unique challenges, and solutions offered for high-resource languages are not necessarily useful for those with fewer resources. Without prioritizing diversity in this emerging field, LLMs will remain vulnerable, with marginalized communities worldwide facing disproportionate risk as LLMs become more widespread.
@@
Computational Constraints
Despite reducing the training data size for each inversion model by 60% to 80%, allowing us to investigate significantly more languages compared to previous studies (Morris et al. 2023; Chen, Lent, and Bjerva 2024), the computational overhead remains substantial. A key limitation of this work is the resource-intensive nature of the experiments, which required approximately 40,000 GPU hours. While expanding this research to include additional languages would further escalate computational costs, we strongly advocate prioritizing NLP security across diverse languages to ensure equitable advancements beyond English.
Defense Mechanisms
This work investigates multilingual inversion attacks in a wide range of languages and exposes patterns of language confusion in these settings. However, the inversion methodology is adopted from Chen, Lent, and Bjerva (2024), and the defense mechanism, the so-called masking defense, proposed in their work also applies to our setting.
LLM’s Cross-lingual Vulnerabilities
Defending multilingual LLMs has proven particularly challenging against attacks such as jailbreak (He et al. 2024), backdoor attacks (Wang et al. 2024a), and fine-tuning (Poppi et al. 2024). We echo these findings by examining multilingual LLMs in embedding inversion attacks across a linguistically diverse set of languages. A major finding of this work is that language confusion can be directly weaponized as an exploit against multilingual LLMs, and as such, we provide further analysis to better understand the underlying patterns of language confusion. Specifically, we present the relations between vulnerabilities and linguistic characteristics to advocate a more linguistically nuanced approach to understanding the vulnerability for linguistically guided mitigation and defense in future research. For example, as indicated in Table 5 and Fig. 2, LLMs trained on language pairs from different scripts and families (specifically for Kazakh and Urdu, or Hindi and Kazakh) are more robust against inversion attacks and language confusion. However, the trade-off between LLM vulnerabilities and LLM performance still needs to be investigated thoroughly in multilingual settings to devise effective defense mechanisms, which we leave for future work.
@@
This work investigates attacks on multilingual embedding models and the phenomenon of language confusion, with the goal of highlighting vulnerabilities across languages representing diverse scripts and typologies. We encourage the community to incorporate a broader range of languages in NLP security research. Some of the languages we examine, such as Yiddish and Meadow Mari, are particularly vulnerable. To minimize potential harm, we do not train any attack models on these languages, instead focusing solely on evaluation to determine their current threat level. The language models used in this study are open-source, ensuring that this research does not pose an immediate threat to EaaS providers, who are likely using private models. Additionally, we take care not to experiment with any sensitive data, ensuring that no real-world harm results from this work.
@@ Acknowledgements
YC, HCL, and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454). We further acknowledge the support of the AAU AI Cloud and express our gratitude to DeiC for providing computing resources on the LUMI cluster (project no. DeiC-AAU-N5-2024085-H2-2024-28). We extend our thanks to Sighvatur Sveinn Davidsson and Frederik Petri Svenningsen for their assistance with AI Cloud, and to Esther Ploeger and Mike Zhang for their discussions related to this work.